Portsmouth Company makes upholstered furniture. Its only variable cost is direct materials. The demand for the company's products far exceeds its manufacturing capacity. The bottleneck (or constriant) in the production process is upholstery labor-hours. Information concerning three of Portsmouth's upholstered chairs appears below:

Recliner

Sofa

Love Seat

Selling price per unit

$

1,092

$

1,900

$

1,420

Variable cost per unit

$

700

$

1,300

$

1,100

Upholstery labor-hours per unit

7 hours

12 hours

4 hours

Required:

1. Portsmouth is considering paying its upholstery laborers additional compensation to work overtime. Assuming that this extra time would be used to produce sofas, up to how much of an overtime premium per hour should the company be willing to pay to keep the upholstery shop open after normal working hours?

2. A small nearby upholstering company has offered to upholster furniture for Portsmouth at a price of $44 per hour. The management of Portsmouth is confident that this upholstering company’s work is high quality and their craftsmen can work as quickly as Portsmouth’s own craftsmen on the simpler upholstering jobs such as the Love Seat. How much additional contribution margin per hour can Portsmouth earn if it provides the raw materials to the nearby company and then hires it to upholster the Love Seats?

3. Should Portsmouth hire the nearby upholstering company?
Answer #1
Concepts and reason

Contribution Margin: It represents the excess of sales over its variable cost. It judges whether the company is able to cover its variable cost and contributes towards the fixed cost.

Fundamentals

Selling Price: Selling price is the price of a product or service at which such product or service is sold to the buyer. Selling price associated with an item is ascertained after rigorous research such that the profitability of the firm is not compromised.

Net Income: The resultant amount after reducing all expenses of the company whether direct or indirect for the period from all revenues is termed as net income.

Direct materials: These are the materials which are used in the Production process which constitutes an integral portion of the final product, the cost of which is identifiable and traceable directly to it. The cost of direct material can be directly chargeable to the final product as compared to the indirect material.

Direct labor: Workers or employees directly involved in the manufacture of goods and services are called direct labor. Direct labor can be directly traceable to a specific product or cost center.

Overhead Cost: It can be referred to as an accounting term which states all the indirect expenses incurred by an organization for operating a business. These expenses are not directly related to the manufacturing of a product or rendering of a service.

Variable Cost: Variable cost is the cost which directly varies with a change in sales. It means to increase/ decrease in sales revenue will have a direct effect on variable cost. There is a linear relation between sales and variable cost. This cost remains fixed per unit but changes in totality. Examples of variable cost are the cost of raw material purchased, direct wages, etc.

Fixed Costs: It is the cost which does not change with increase or decrease in the production output level. It remains constant throughout the production. It is a part of the total cost to run a business along with variable cost.

1.

Compute the contribution per labor hour for the sofa, using MS Excel as follows:

Computation of Contribution per labor hour for Sofa
Particulars
Amount
Selling Price per unit
$1,900
Less: Variable Price per

Hence, the company should be willing to pay a maximum of $50 as an overtime premium.

Working Notes:

Compute the contribution margin per unit using the equation as follows:

Contribution=SellingpriceVariablecost=$1,900$1,300=$600\begin{array}{c}\\{\rm{Contribution}} = {\rm{Selling price}} - {\rm{Variable cost}}\\\\ = \$ 1,900 - \$ 1,300\\\\ = \$ 600\\\end{array}

Hence, the contribution per unit is $600.

Compute the contribution per labor hour using the equation as follows:

Contributionperlaborhour=ContributionmarginLaborhoursrequired=$60012=$50\begin{array}{c}\\{\rm{Contribution per labor hour}} = \frac{{{\rm{Contribution margin}}}}{{{\rm{Labor hours required}}}}\\\\ = \frac{{\$ 600}}{{12}}\\\\ = \$ 50\\\end{array}

Hence, the contribution per labor hour for sofa is $50.

2.

Compute the contribution per labor hour for recliner using MS Excel as follows:

Computation of Contribution per labor hour for Recliner
Particulars
Amount
Selling Price per unit
$1,092
Less: Variable Price

Working Notes:

Compute the contribution margin per unit using the equation as follows:

Contribution=SellingpriceVariablecost=$1,092$700=$392\begin{array}{c}\\{\rm{Contribution}} = {\rm{Selling price}} - {\rm{Variable cost}}\\\\ = \$ 1,092 - \$ 700\\\\ = \$ 392\\\end{array}

Hence, the contribution per unit is $392.

Compute the contribution per labor hour using the equation as follows:

Contributionperlaborhour=ContributionmarginLaborhoursrequired=$3927=$56\begin{array}{c}\\{\rm{Contribution per labor hour}} = \frac{{{\rm{Contribution margin}}}}{{{\rm{Labor hours required}}}}\\\\ = \frac{{\$ 392}}{7}\\\\ = \$ 56\\\end{array}

Hence, the contribution per labor hour for recliner is $56.

Compute the contribution per labor hour of love seat using MS-Excel as follows:

Computation of Contribution per labor hour for Love Seat
Particulars
Amount
Selling Price per unit
$1,420
Less: Variable Pric

Hence, the additional contribution per labor hour earned is $36.

Working Notes:

Compute the contribution margin per unit using the equation as follows:

Contribution=SellingpriceVariablecost=$1,420$1,100=$320\begin{array}{c}\\{\rm{Contribution}} = {\rm{Selling price}} - {\rm{Variable cost}}\\\\ = \$ 1,420 - \$ 1,100\\\\ = \$ 320\\\end{array}

Hence, the contribution per unit is $320.

Compute the contribution per labor hour using the equation as follows:

Contributionperlaborhour=ContributionmarginLaborhoursrequired=$3204=$80\begin{array}{c}\\{\rm{Contribution per labor hour}} = \frac{{{\rm{Contribution margin}}}}{{{\rm{Labor hours required}}}}\\\\ = \frac{{\$ 320}}{4}\\\\ = \$ 80\\\end{array}

Hence, the contribution per labor hour for love seat is $80.

Compute the additional contribution margin earned using the equation as follows:

Additionalmargin=ContributionmarginperlaborhourPricechargedbyanotherfirm=$80$44=$36\begin{array}{c}\\{\rm{Additional margin}} = {\rm{Contribution margin per labor hour}} - {\rm{Price charged by another firm}}\\\\ = \$ 80 - \$ 44\\\\ = \$ 36{\rm{ }}\\\end{array}

Hence, the additional contribution per labor hour earned is $36.

3.

The contribution margin per labor hour of all the three furniture products is more than the price charged by the upholstering company that is $44. Hence, the upholstering company should be hired. The freed up time could be used to produce more love seats as the contribution per labor hour is highest for love seats.

Ans: Part 1

The company should be willing to pay a maximum $50 per hour as overtime premium.